OpenStudy (anonymous):

Which of the following is equivalent to the complex number i^7? a.1 b. i c.-1 d.- i

OpenStudy (anthonyym):

Do you know what i is?

OpenStudy (mathmale):

I found this question confusing until I realized that you meant i^7, or \[i^7.\] That makes a BIG difference. Note that i^7=i^4*i^3. What is i^3 equivalent to?

OpenStudy (anonymous):

yeah sorry I went back and changed it.

OpenStudy (mathmale):

While we're on the topic: i is the imaginary operator, and is the definition of sqrt(-1). Note the following:\[i^1=i;~ i^2=-1;~ i^3=-i; ~and~i^4=1. \]

OpenStudy (anonymous):

thanks that is really helpful

OpenStudy (mathmale):

This pattern repeats itself. So, if you start with i^7, you can replace that with i^3. What is i^3?

OpenStudy (anonymous):

-i

OpenStudy (mathmale):

Yes. so now you can choose the correct answer choice. nice work!

OpenStudy (anonymous):

thank you so much, I'll give you a medal!!

OpenStudy (mathmale):

Thanks, miiles! :)